OpenStudy (anonymous):

Find a cubic function with the given zeros. √2, -√2 , -2

OpenStudy (fifciol):

\[(x+2)(x-\sqrt2)(x+\sqrt2)=(x+2)(x^2-2)=x^3-2x+2x^2-4=x^3+2x^2-2x-4\]

OpenStudy (anonymous):

Thanks~